OpenStudy (anonymous):

Find the solution to the system of equations. 3x-3y-z=-7 -x+3y-z=3 x + 4z=18

OpenStudy (anonymous):

A. x=6, y=0, z=8 B. x=3, y=0, z=2 C. x=2, y=4, z=3 D. x=2, y=3, z=4

OpenStudy (solomonzelman):

You can choose to set a matrix, or make the following substitution, using your 3rd equation. \(\large\color{black}{ \displaystyle x + 4z=18 \quad \Longrightarrow\quad \color{#558822}{x=18-4z}}\) And then you would get this (below), in the remaining two (1st and 2nd) equations. \(\large\color{black}{ \displaystyle 3\left(\color{#558822}{18-4z}\right)-3y-z=-7 }\) \(\large\color{black}{ \displaystyle -\left(\color{#558822}{18-4z}\right)+3y-z=3 }\) By doing this you eliminate one parameter, and thus your equation is easier to solve.